OpenStudy (anonymous):

what is 2m^3n^-6/10m^2n^3 ? help plz

OpenStudy (mathstudent55):

Is this it? \(\large \dfrac{2m^3n^{-6}}{10m^2n^3}\)

OpenStudy (anonymous):

Yeah that is it.

OpenStudy (mathstudent55):

Ok. How do you divide powers? What is \(\large \dfrac{5^7}{5^3} =\)

OpenStudy (anonymous):

[5^{4}\]

OpenStudy (mathstudent55):

Correct. When you divide powers with the same base, you subtract the exponents.

OpenStudy (mathstudent55):

Now think of this problem as three separate problems: \(\large \dfrac{2}{10} \times \dfrac{m^3}{m^2} \times \dfrac{n^{-6}}{n^3} \) Simplify each fraction and multiply them together.

OpenStudy (anonymous):

Is it m/5n^3?

OpenStudy (anonymous):

no its not..

OpenStudy (anonymous):

it should be m/5n^9

OpenStudy (mathstudent55):

\(\large \dfrac{2}{10} \times \dfrac{m^3}{m^2} \times \dfrac{n^{-6}}{n^3}\) \( = \large \dfrac{1}{5} \times m^{3-2} \times n^{-6-3}\) \( = \large \dfrac{1}{5} \times m^1 \times n^{-9}\) \(= \large \dfrac{m}{5n^9}\)